# Optimal Software Pipelining using an SMT-Solver

### v1

@@ 1.3 Loop Unrolling
A straightforward way of optimizing this loop is to apply loop unrolling. Unrolling the loop four times yields the schedule in Figure 2(a). As illustrated, when loop unrolling, the body of a loop is replicated multiple times. In this case, the body of the newly created loop contains instructions from four iterations (in the picture, each iteration is given a different color). This creates more opportunity for parallelism, improving the efficiency of the execution. Executing the unrolled loop takes $256\times 7=1792$ cycles, a speed-up of a factor 2.29. The drawback of unrolling is the increased code size. In this case, the code size has increased by a factor $1.75$ to $7$ cycles.
@@ 2.1 Modulo Scheduling
    However, when modulo scheduling, we also need to make sure that operation $b$ reads the value written by operation $a$ of the current iteration before operation $a$ of the next iteration overwrites it with a potentially different value. To do so, we introduce a loop-carried dependency between $b$ and $a$ with latency 0 and distance $1$. Here, the distance indicates the number of iterations $b$ and $a$ are apart. So, the dependency means that operation $b$ from iteration $n$ should be scheduled before operation $a$ of iteration $n+1$. In the modulo schedule for a specific $\mathrm{II}$, each $\mathrm{II}$ cycles a new iteration is started. So, to meet this dependency constraint $b$ should be scheduled at most $\mathrm{II}$ cycles after $a$. In general, to meet a dependency constraint between $o_{1}$ and $o_{2}$ with latency $l$ and distance $d$, the following should hold[12]:
@@
    For each cycle $c$ with $0\leq c\leq\mathrm{MaxCycle}$, issue-slot output port $op$, and bus $b$, a Boolean variable $\mathsf{Connected}_{c,op,b}$ modeling whether at cycle $c$ the connection between issue-slot output-port $op$ and bus $b$ is made.
@@
    For each cycle $c$ with $0\leq c\leq\mathrm{MaxCycle}$, bus $b$, and register-file write port $wp$, a Boolean variable $\mathsf{Connected}_{c,b,wp}$ modeling whether at cycle $c$ the connection between bus $b$ and register-file write-port $wp$ is made.
@@
    For each data flow dependency between operations $o_{1}$ and $o_{2}$, issue-slots $s_{1}$ and $s_{2}$, cycle $c$, let $op$ be the issue-slot outport that $o_{1}$ writes the data to when it is scheduled on slot $s_{1}$, let $W$ be the set of register-file write-ports of the register-file that $o_{2}$ reads the data from when it is scheduled on $s_{2}$, and let $\mathrm{B}(wp,op)$ be the set of buses that connect register-file write-port $wp$ with issue-slot output-port opop
@@
    For all cycles $c_{1}$ and $c_{2}$ such that $c_{1}\%\mathrm{II}=c_{2}\%\mathrm{II}$, register-file write port $wp$, and distinct buses $b_{1}$ and $b_{2}$:
@@ 2.5 Extension of the SMT Problem to deal with Register Pressure
    For each virtual register $v$ and register-file $RF$, integer variables
@@
    • $\mathrm{End}_{RF,v}$ modeling the cycle in which $v$ ends becoming live in register-file $RF$.
    • $\mathrm{MinUse}_{RF,v}$ modeling the cycle in which $v$ is used first in register-file $RF$.
    • $\mathrm{MaxUse}_{RF,v}$ modeling the cycle in which $v$ is used last in register-file $RF$.
    • $\mathrm{MaxUseBeforeDef}_{RF,v}$ modeling the cycle in which $v$ is used last before it is defined in register-file $RF$.
@@
    For each virtual register $v$ and register-file $RF$, Boolean variables
    • $\mathrm{Used}_{RF,v}$ modeling whether $v$ is used in register-file $RF$.
    • $\mathrm{UsedBeforeDef}_{RF,v}$ modeling whether $v$ is used before it is defined in register-file $RF$.
    • $\mathrm{LiveOut}_{RF,v}$ modeling whether $v$ is live-out in register-file $RF$.
@@
    For each virtual register $v$, register-file $RF$, and cycle $c$, Boolean variables
    • $\mathrm{Live}_{RF,v,c}$ modeling whether $v$ is live in register-file $RF$ at cycle $c$
@@
    Define $\mathrm{Used}$, $\mathrm{MinUse}$, $MaxUse$, and $\mathrm{MaxUseBeforeDef}$ in terms of $\mathrm{Cycle}$ and $\mathrm{Slot}$
    For each virtual register $v$, let $U$ be the set of operations in the loop that uses $v$, and for each register-file $RF$, let $S_{RF,u,v}$ be the set of issue-slots such that $v$ is read from register-file $RF$ if $u$ is scheduled on the slot.
@@
    For each virtual register $v$, let $U$ be the set of operations outside, but reachable from, the loop that use $v$, and for each register-file $RF$, let $S_{RF,u,v}$ be the set of issue-slots such that operation $u$ reads virtual register $v$ from register-file $RF$ if $u$ is scheduled on the slot.
@@
    For each virtual register $v$, register-file $RF$, and cycle $c$ such that $0\leq c<\mathrm{II}$:
@@
    For each register-file $RF$ with capacity capcap and cycle $c$ such that $0\leq c<\mathrm{II}$:

### v2

@@ 1.3 Loop Unrolling
A straightforward way of optimizing this loop is to apply loop unrolling. Unrolling the loop four times yields the schedule in Figure 2(a). As illustrated, when loop unrolling, the body of a loop is replicated multiple times. In this case, the body of the newly created loop contains instructions from four iterations (in the picture, each iteration is given a different color). This creates more opportunity for parallelism, improving the efficiency of the execution. Executing the unrolled loop takes $256\times 7=1792$ cycles, a speed-up of a factor 2.29. The drawback of unrolling is the increased code size. In this case, the code size has increased by a factor $1.75$ to 77 cycles.
@@ 2.1 Modulo Scheduling
    However, when modulo scheduling, we also need to make sure that operation $b$ reads the value written by operation $a$ of the current iteration before operation $a$ of the next iteration overwrites it with a potentially different value. To do so, we introduce a loop-carried dependency between $b$ and $a$ with latency 0 and distance 11. Here, the distance indicates the number of iterations $b$ and $a$ are apart. So, the dependency means that operation $b$ from iteration $n$ should be scheduled before operation $a$ of iteration $n+1$. In the modulo schedule for a specific $\mathrm{II}$, each $\mathrm{II}$ cycles a new iteration is started. So, to meet this dependency constraint $b$ should be scheduled at most $\mathrm{II}$ cycles after $a$. In general, to meet a dependency constraint between $o_{1}$ and $o_{2}$ with latency $l$ and distance $d$, the following should hold[12]:
@@
    For each cycle $c$ with $0\leq c\leq\mathrm{MaxCycle}$, issue-slot output port opop, and bus $b$, a Boolean variable $\mathsf{Connected}_{c,op,b}$ modeling whether at cycle $c$ the connection between issue-slot output-port opop and bus $b$ is made.
@@
    For each cycle $c$ with $0\leq c\leq\mathrm{MaxCycle}$, bus $b$, and register-file write port wpwp, a Boolean variable $\mathsf{Connected}_{c,b,wp}$ modeling whether at cycle $c$ the connection between bus $b$ and register-file write-port wpwp is made.
@@
    For each data flow dependency between operations $o_{1}$ and $o_{2}$, issue-slots $s_{1}$ and $s_{2}$, cycle $c$, let opop be the issue-slot outport that $o_{1}$ writes the data to when it is scheduled on slot $s_{1}$, let $W$ be the set of register-file write-ports of the register-file that $o_{2}$ reads the data from when it is scheduled on $s_{2}$, and let $\mathrm{B}(wp,op)$ be the set of buses that connect register-file write-port wpwp with issue-slot output-port opop
@@
    For all cycles $c_{1}$ and $c_{2}$ such that $c_{1}\%\mathrm{II}=c_{2}\%\mathrm{II}$, register-file write port wpwp, and distinct buses $b_{1}$ and $b_{2}$:
@@ 2.5 Extension of the SMT Problem to deal with Register Pressure
    For each virtual register $v$ and register-file RFRF, integer variables
@@
    • $\mathrm{End}_{RF,v}$ modeling the cycle in which $v$ ends becoming live in register-file RFRF.
    • $\mathrm{MinUse}_{RF,v}$ modeling the cycle in which $v$ is used first in register-file RFRF.
    • $\mathrm{MaxUse}_{RF,v}$ modeling the cycle in which $v$ is used last in register-file RFRF.
    • $\mathrm{MaxUseBeforeDef}_{RF,v}$ modeling the cycle in which $v$ is used last before it is defined in register-file RFRF.
@@
    For each virtual register $v$ and register-file RFRF, Boolean variables
    • $\mathrm{Used}_{RF,v}$ modeling whether $v$ is used in register-file RFRF.
    • $\mathrm{UsedBeforeDef}_{RF,v}$ modeling whether $v$ is used before it is defined in register-file RFRF.
    • $\mathrm{LiveOut}_{RF,v}$ modeling whether $v$ is live-out in register-file RFRF.
@@
    For each virtual register $v$, register-file RFRF, and cycle $c$, Boolean variables
    • $\mathrm{Live}_{RF,v,c}$ modeling whether $v$ is live in register-file RFRF at cycle $c$
@@
    Define $\mathrm{Used}$, $\mathrm{MinUse}$, MaxUseMaxUse, and $\mathrm{MaxUseBeforeDef}$ in terms of $\mathrm{Cycle}$ and $\mathrm{Slot}$
    For each virtual register $v$, let $U$ be the set of operations in the loop that uses $v$, and for each register-file RFRF, let $S_{RF,u,v}$ be the set of issue-slots such that $v$ is read from register-file RFRF if $u$ is scheduled on the slot.
@@
    For each virtual register $v$, let $U$ be the set of operations outside, but reachable from, the loop that use $v$, and for each register-file RFRF, let $S_{RF,u,v}$ be the set of issue-slots such that operation $u$ reads virtual register $v$ from register-file RFRF if $u$ is scheduled on the slot.
@@
    For each virtual register $v$, register-file RFRF, and cycle $c$ such that $0\leq c<\mathrm{II}$:
@@
    For each register-file RFRF with capacity capcap and cycle $c$ such that $0\leq c<\mathrm{II}$: